# supplement

### v1

@@ 1 Introduction
In this paper, we derive a novel non-asymptotic uncertainty bound for kernel-based estimation assuming a general bound on the noise energy. In particular, we derive an analytical solution that exactly characterizes the worst-case realization within the function hypothesis class. This uncertainty bound has the same structure as the high-probability uncertainty bounds from GP regression [Srinivas et al., 2012; Abbasi-Yadkori, 2013; Fiedler et al., 2021; Molodchyk et al., 2025], but with a measurement noise covariance $\sigma^{2}$ that depends on the test input location. Furthermore, we show that the derived bound recovers results from kernel interpolation [Weinberger and Golomb, 1959; Wendland, 2004] and linear regression [Fogel, 1979] as special cases. Finally, we contrast the proposed robust treatment to existing bounds for GP regression.
@@
The matrix $I_{n}\in\mathbb{R}^{n\times n}$ denotes the identity matrix of dimension $n$. For a symmetric positive-semidefinite matrix $A\in\mathbb{R}^{n\times n}$, $A^{1/2}$ denotes the (positive-semidefinite) symmetric matrix square root, i.e., $A^{1/2}A^{1/2}=A$, and $\|x\|_{A}^{2}\doteq x^{\top}Ax$ denotes the weighted Euclidean norm of a vector $x\in\mathbb{R}^{n}$. The Dirac delta function is denoted by $\delta:\mathbb{R}^{n_{x}}\rightarrow\mathbb{R}$, with $\delta(x)=1$ for $x=0$ and $\delta(x)=0$ otherwise. Superscripts $f$ and $w$ will refer to the latent function and the noise, respectively. Accordingly, the kernel function is denoted by $k^{\square}\colon\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{x}}\to\mathbb{R}$, with $\square\in\{f,w\}$, and the associated RKHS is denoted by $\mathcal{H}_{k^{\square}}$. For two arbitrary ordered sets of indices $\mathbb{I},\mathbb{J}\subseteq\mathbb{N}$, the matrix $K^{\square}_{\mathbb{I},\mathbb{J}}\doteq[k^{\square}(x_{i},x_{j})]_{i\in\mathbb{I},j\in\mathbb{J}}$ is the Gram matrix collecting the evaluations of the kernel function $k^{\square}$ at pairs of input locations $x_{i}$, $x_{j}$, with $i\in\mathbb{I}$ and $j\in\mathbb{J}$. We denote by $1:N\doteq\{1,\dots,N\}\subseteq\mathbb{N}$ the set of indices for the training data-points, while we use $x_{N+1}$ to represent the arbitrary test input. For instance, $K^{f}_{N+1,1:N}\in\mathbb{R}^{1\times N}$ corresponds to $\begin{bmatrix}k^{f}(x_{N+1},x_{1}),\ldots,k^{f}(x_{N+1},x_{N})\end{bmatrix}$ and can be interpreted as the covariance matrix between the test- and training input locations.
@@ 2 Problem set-up
We consider the problem of estimating an unknown latent function $f^{\mathrm{tr}}\colon\mathcal{X}\to\mathbb{R}$, with $\mathcal{X}\subseteq\mathbb{R}^{n_{x}}$, from noisy measurements
@@
collected at known training input locations $x_{i}\in\mathcal{X}$. Our goal is to compute worst-case uncertainty bounds around the latent function $f^{\mathrm{tr}}$ given the observed data set $\mathcal{D}\doteq\{(x_{i},y_{i})\}_{i=1}^{N}$, which is subject to the unknown noise $w^{\mathrm{tr}}\colon\mathcal{X}\to\mathbb{R}$. We phrase the problem in the framework of estimation in RKHSs [Aronszajn, 1950; Berlinet and Thomas-Agnan, 2004], and we model both the latent function $f^{\mathrm{tr}}$ and the noise $w^{\mathrm{tr}}$ as elements of a RKHS with a known kernel and a bound on their RKHS norm.
@@
Characterizing boundedness of the noise using a noise kernel $k^{w}$ and an RKHS-norm bound $\Gamma_{w}^{2}$ provides a very general description and can model various scenarios, such as energy-bounded noise – see Section 4.1 for more details. Notably, we require no assumption regarding the distribution or independence of the noise, the latent function, or the input locations.
@@ 3 Kernel regression bounds for energy-bounded noise
In the following, we present the main result of our paper, determining point-wise and tight uncertainty bounds $\underline{f}(x_{N+1})\leq f^{\mathrm{tr}}(x_{N+1})\leq\overline{f}(x_{N+1})$ for the value of the latent function at an arbitrary test point $x_{N+1}\in\mathcal{X}$. This task can be formulated as an infinite-dimensional optimization problem, taking the bounded-RKHS-norm assumption into account. The optimal upper bound $\overline{f}(x_{N+1})$ is defined as
$\overline{f}(x_{N+1})=\sup_{\begin{subarray}{c}f\in\mathcal{H}_{k^{f}},\\ w\in\mathcal{H}_{k^{w}},\\ \end{subarray}}\quad$ $f(x_{N+1})$ (2a)
$\mathrm{s.t.}\quad$ $f(x_{i})+w(x_{i})=y_{i},\>i=1,\ldots,N$ (2b)
$\|f\|_{\mathcal{H}_{k^{f}}}^{2}\leq\Gamma_{f}^{2}$ (2c)
$\|w\|_{\mathcal{H}_{k^{w}}}^{2}\leq\Gamma_{w}^{2}$ (2d)
@@
$\underline{f}(x_{N+1})=\inf_{\begin{subarray}{c}f\in\mathcal{H}_{k^{f}},\\ w\in\mathcal{H}_{k^{w}},\\ \end{subarray}}\quad$ $f(x_{N+1})$ (3a)
$\mathrm{s.t.}\quad$ (2b)-(2d) (3b)
In the following, we focus on computing the upper bound $\overline{f}(x_{N+1})$; for the lower bound, the presented results in Sections 3.1 and 3.2 are analogously derived in Appendices B and C, respectively.
Stated as in (2), the optimization problem is infinite-dimensional and is not directly tractable. Our key result, presented in the remainder of the section, consists in deriving an analytical solution for the bound evaluated at an arbitrary input location. To this end, we first study a relaxed formulation of this optimization problem in Section 3.1. In Section 3.2, we discuss how to recover the optimal solution from the relaxed problem.
@@ 3.1 Relaxed solution
The relaxed formulation of optimization problem (2) considers the sum of the RKHS-norm constraints (2c), (2c) instead of enforcing them individually:
$\overline{f}^{\sigma}(x_{N+1})=\sup_{\begin{subarray}{c}f\in\mathcal{H}_{k^{f}},\\ w\in\mathcal{H}_{k^{\sigma}},\\ \end{subarray}}\quad$ $f(x_{N+1})$ (4a)
$\mathrm{s.t.}\quad$ $f(x_{i})+w(x_{i})=y_{i},\>i=1,\ldots,N$ (4b)
@@
defines the maximum norm (4c) in the RKHS $\mathcal{H}_{k^{f}+k^{\sigma}}$ based on Assumption 1, reduced by the minimum norm required to interpolate the data.
@@
Let Assumptions 1 and 2 hold. Then, the solution of Problem (4) is given by
$\overline{f}^{\sigma}(x_{N+1})=f^{\mu}_{\sigma}+\beta_{\sigma}\sqrt{\Sigma^{f}_{\sigma}(x_{N+1})}.$ (8)
Sketch of proof: First, following arguments from the Representer Theorem [Kimeldorf and Wahba, 1971; Schölkopf et al., 2001], we show that the solution of Problem (4) is finite-dimensional. Next, two coordinate transformations are employed to reduce the number of free variables resulting from the interpolation constraint (4b), and to address the possible rank-deficiency of the kernel matrix $K^{f}_{1:N+1,1:N+1}$ for the latent function at the test and training input locations. Finally, the problem is reduced to an equivalent linear program with a norm-ball constraint that can be analytically solved. Expressing the solution in terms of the original coordinates then leads to (8). The detailed proof can be found in Appendix B.
@@ 3.2 Optimal solution
Let Assumptions 1 and 2 hold. Then, the solution of Problem (2) is given by
@@
Sketch of proof: Similarly to Lemma 1, we first show that Problem (2) admits a finite-dimensional representation. The latter is analyzed depending on which of the constraints (2c) and (2d) are active, i.e., influence the optimal solution and have a corresponding strictly positive optimal Lagrange multiplier $\lambda^{f,\star},\lambda^{w,\star}$. This leads to three scenarios: In Case 1, it holds that $\lambda^{f,\star}>0$, $\lambda^{w,\star}=0$ and the optimal solution of (2) can be recovered by the relaxed problem for $\sigma^{\star}\rightarrow\infty$, for which the combined RKHS-norm constraint (4c) reduces to (2c). In Case 2, $\lambda^{f,\star}=0$, $\lambda^{w,\star}>0$ and $\sigma^{\star}\rightarrow 0$ recovers the optimal solution, rendering the constraint (4c) equivalent to (2d). In Case 3 both constraints are active and the optimal noise parameter is determined by $\sigma^{\star}=\sqrt{\lambda^{f,\star}/\lambda^{w,\star}}\in(0,\infty)$, i.e., the ratio of the optimal Lagrange multipliers. The set of active constraints at the optimal solution can be determined by case distinction, based on feasibility of the primal solutions for the neglected constraints. Finally, it is shown that the optimal noise parameter $\sigma^{\star}$ in all three cases minimizes (9). This is illustrated in Fig. 1, which depicts the optimal noise parameters $\sigma^{\star}_{\sup}$$\sigma^{\star}_{\inf}$, for which the relaxed upper and lower bound, $\overline{f}^{\sigma}(x_{N+1})$ and $\underline{f}^{\sigma}(x_{N+1})$, correspond to the optimal bounds, $\overline{f}(x_{N+1})$ and $\underline{f}(x_{N+1})$, respectively. The detailed proof can be found in Appendix C.
@@
Figure 1: Illustrative example for Theorem 1. The optimal upper and lower bounds (solid black) for the (unknown) latent function $f^{\mathrm{tr}}$ (dashed white) are determined by the relaxed bounds (shaded) around the GP posterior mean (dotted black) for an optimal choice of noise parameter $\sigma^{\star}_{\sup}$ (upper bound) and $\sigma^{\star}_{\inf}$ (lower bound). The three upper plots show the relaxed upper and lower bounds, $\overline{f}^{\sigma}$ and $\underline{f}^{\sigma}$ for the values $\sigma=\{10^{2},10^{0},10^{-2}\}$, respectively. The two bottom colorbars indicate the respective optimal values $\sigma^{\star}_{\sup}$ and $\sigma^{\star}_{\inf}$ for the upper and lower bound. The plotted relaxed upper (lower) bounds equal the optimal upper (lower) bound for each test point where the color of the shaded area matches the color indicated in the colorbar for the optimal value $\sigma^{\star}_{\sup}$ ($\sigma^{\star}_{\inf}$).
@@
Theorem 1 reduces the solution of the infinite-dimensional optimization problem (2) to a scalar, unconstrained optimization problem over the noise parameter $\sigma$. As such, it is amenable for efficient iterative optimization. Since running a fixed number of iterations of, e.g., gradient descent applied to Problem (9) returns a valid, improved, upper bound $\overline{f}^{\sigma}(x_{N+1})$, this allows for iterative refinement of the uncertainty envelope. The solution thereby obtained can thus be easily integrated into existing pipelines for downstream tasks, such as uncertainty quantification in streaming-data settings or model-based reinforcement learning [Deisenroth and Rasmussen, 2011; Berkenkamp et al., 2017; Kamthe and Deisenroth, 2018].
@@ 3.3 Special cases
For both cases with only one active constraint, the optimal bound can be determined directly in closed form, without optimizing for the noise parameter $\sigma$. In the following, we provide the respective optimal solutions, as well as easy-to-evaluate expressions for determining the active constraint set. Noteworthy, the analytic solutions recover known bounds in specific regression settings, highlighting that the proposed bound is a generalization thereof; we detail these connections in Section 4.1.
@@
Let Assumptions 1 and 2 hold. Define $P\doteq(\Phi_{1:N}^{\top}\left(K^{w}_{1:N,1:N}\right)^{-1}\Phi_{1:N})^{-1}$ and $\theta^{\mu}\doteq P\Phi_{1:N}^{\top}\left(K^{w}_{1:N,1:N}\right)^{-1}y$. Then, if
@@
Case 2 can happen in two scenarios: For finite-dimensional hypothesis spaces, i.e., $f^{\mathrm{tr}}(\cdot)=\begin{bmatrix}\phi_{1}(\cdot)&\ldots&\phi_{r}(\cdot)\end{bmatrix}\theta^{\mathrm{tr}}$ for some features $\phi_{i}(\cdot)$, $i=1,\ldots,r$, the latent function $f^{\star}(x_{N+1})$ generally does not have sufficient degrees of freedom to interpolate an arbitrary data set. As such, the optimal bound $\overline{f}(x_{N+1})$ in (13) consists of two components, the value of the least-squares estimator $f^{\mu}(x_{N+1})\doteq\Phi_{N+1}\theta^{\mu}$, as well as a term proportional to the maximum RKHS norm of the noise $\Gamma_{w}^{2}$, subtracted by $y^{\top}(K^{w}_{1:N,1:N})^{-1}y-\|\theta^{\mu}\|_{P^{-1}}$, the minimum RKHS norm required to eliminate the offset between the least-squares estimator and the data. For infinite-dimensional hypothesis spaces, the latent function $f^{\star}$ can generally interpolate the offset between the optimal noise function $w^{\star}$ and the training data; however, neglecting the RKHS-norm constraint (2c) on $f^{\star}$ only leads to sensible estimates when the test point coincides with a training input location. In this case, the feature vector $\Phi_{N+1}\in\mathbb{R}^{(N+1)\times r}$ has rank $N$, which simplifies the general result in Proposition 2.
@@
Let Assumptions 1 and 2 hold. Suppose that $K^{f}_{1:N,1:N}$ is invertible and $x_{N+1}=x_{k}\in\mathbb{X}$. Then, if
$\left\|\theta^{\star}\right\|_{2}^{2}$ $=\left\|y+K^{w}_{1:N,k}\frac{\Gamma_{w}}{\sqrt{K^{w}_{k,k}}}\right\|_{(K^{f}_{1:N,1:N})^{-1}}^{2}\leq\Gamma_{f}^{2},$ (14)

### v3

# Optimal kernel regression bounds under energy-bounded noise

@@ 1 Introduction
In this paper, we propose a novel non-asymptotic uncertainty bound for kernel-based estimation assuming a general bound on the noise energy. In particular, for each query input location, the proposed bound exactly characterizes the worst-case latent function within the given hypothesis class. The obtained uncertainty bound has the same structure as the high-probability uncertainty bounds from GP regression [Srinivas et al., 2012; Abbasi-Yadkori, 2013; Fiedler et al., 2021; Molodchyk et al., 2025], but with a measurement noise covariance $\sigma^{2}$ that depends on the test input location. Furthermore, we show that the derived bound recovers results from kernel interpolation [Weinberger and Golomb, 1959; Wendland, 2004] and linear regression [Fogel, 1979] as special cases. Finally, we contrast the proposed robust treatment to existing bounds for GP regression.
@@
The matrix $I_{n}\in\mathbb{R}^{n\times n}$ denotes the identity matrix of dimension $n$. For a symmetric positive-semidefinite matrix $A\in\mathbb{R}^{n\times n}$, $A^{1/2}$ denotes the (positive-semidefinite) symmetric matrix square root, i.e., $A^{1/2}A^{1/2}=A$, and $\|x\|_{A}^{2}\doteq x^{\top}Ax$ denotes the weighted Euclidean norm of a vector $x\in\mathbb{R}^{n}$. The Dirac delta function is denoted by $\delta:\mathbb{R}^{n_{x}}\rightarrow\mathbb{R}$, with $\delta(x)=1$ for $x=0$ and $\delta(x)=0$ otherwise. Superscripts $f$ and $w$ will refer to the latent function and the noise, respectively. Accordingly, the kernel function is denoted by $k^{\square}\colon\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{x}}\to\mathbb{R}$, with $\square\in\{f,w\}$, and the associated RKHS is denoted by $\mathcal{H}_{k^{\square}}$. For two arbitrary ordered sets of indices $\mathbb{I},\mathbb{J}\subseteq\mathbb{N}$, the matrix $K^{\square}_{\mathbb{I},\mathbb{J}}\doteq[k^{\square}(x_{i},x_{j})]_{i\in\mathbb{I},j\in\mathbb{J}}$ is the Gram matrix collecting the evaluations of the kernel function $k^{\square}$ at pairs of input locations $x_{i}$, $x_{j}$, with $i\in\mathbb{I}$ and $j\in\mathbb{J}$. We denote by $1:N\doteq\{1,\dots,N\}\subseteq\mathbb{N}$ the set of indices for the training data points, while we use $x_{N+1}$ to represent the arbitrary test input. For instance, $K^{f}_{N+1,1:N}\in\mathbb{R}^{1\times N}$ corresponds to $\begin{bmatrix}k^{f}(x_{N+1},x_{1}),\ldots,k^{f}(x_{N+1},x_{N})\end{bmatrix}$ and can be interpreted as the covariance matrix between the test- and training-input locations.
@@ 2 Problem set-up
We consider the problem of estimating an unknown latent function $f^{\mathrm{tr}}\colon\mathcal{X}\to\mathbb{R}$, with $\mathcal{X}\subseteq\mathbb{R}^{n_{x}}$ from noisy measurements
@@
collected at known training input locations $x_{i}\in\mathcal{X}$. Our goal is to compute worst-case uncertainty bounds around the latent function $f^{\mathrm{tr}}$ given the observed data set $\mathcal{D}\doteq\{(x_{i},y_{i})\}_{i=1}^{N}$, which is subject to the unknown noise $w^{\mathrm{tr}}\colon\mathcal{X}\to\mathbb{R}$. We phrase the problem in the framework of estimation in RKHSs [Aronszajn, 1950; Berlinet and Thomas-Agnan, 2004], and we model both the latent function $f^{\mathrm{tr}}$ and the noise $w^{\mathrm{tr}}$ as elements of an RKHS with a known kernel and a bound on their RKHS norm.
@@
Characterizing boundedness of the noise using a kernel $k^{w}$ and an RKHS-norm bound $\Gamma_{w}^{2}$ provides a very general description and can model various scenarios: for instance, it captures the setting in which the noise sequence has bounded energy, as we elucidate in Section˜4.1. Additionally, modeling noise as a deterministic quantity allows us to by-pass additional assumptions on the distribution or independence of the noise, latent function or input locations.
@@ 3 Kernel regression bounds for energy-bounded noise
In the following, we present the main result of our paper, determining tight point-wise uncertainty bounds $\underline{f}(x_{N+1})\leq f^{\mathrm{tr}}(x_{N+1})\leq\overline{f}(x_{N+1})$ for the value of the latent function at an arbitrary test point $x_{N+1}\in\mathcal{X}$. This task can be formulated as an infinite-dimensional optimization problem, taking the bounded-RKHS-norm assumption into account. The optimal upper bound $\overline{f}(x_{N+1})$ is defined as
$\overline{f}(x_{N+1})=\sup_{\begin{subarray}{c}f\in\mathcal{H}_{k^{f}},\\ w\in\mathcal{H}_{k^{w}}\\ \end{subarray}}\quad$ $f(x_{N+1})$ (2a)
$\mathrm{s.t.}\quad$ $f(x_{i})+w(x_{i})=y_{i},\>i=1,\ldots,N,$ (2b)
$\|f\|_{\mathcal{H}_{k^{f}}}^{2}\leq\Gamma_{f}^{2},$ (2c)
$\|w\|_{\mathcal{H}_{k^{w}}}^{2}\leq\Gamma_{w}^{2}.$ (2d)
@@
$\underline{f}(x_{N+1})=\inf_{\begin{subarray}{c}f\in\mathcal{H}_{k^{f}},\\ w\in\mathcal{H}_{k^{w}}\\ \end{subarray}}\quad$ $f(x_{N+1})$ (3a)
[…]
In the following, we focus on computing the upper bound $\overline{f}(x_{N+1})$; for the lower bound, the presented results in Sections˜3.1 and 3.2 are analogously derived in Appendices˜B and C, respectively.
Stated as in (2), the optimization problem is infinite-dimensional and is not directly tractable. Our key result, presented in the remainder of the section, consists in finding an exact reformulation of the constrained, infinite-dimensional problem (2) as a scalar, unconstrained one. The solution of the latter at an arbitrary input location is expressed in terms of familiar quantities from Gaussian process regression, for an optimal choice of measurement noise covariance. We present our derivation by first studying a relaxed formulation of this optimization problem in Section˜3.1. Then, in Section˜3.2, we discuss how to recover the optimal solution from the relaxed problem.
@@ 3.1 Relaxed solution
The relaxed formulation of optimization problem (2) considers the sum of the RKHS-norm constraints (2c), (2d) instead of enforcing them individually:
$\overline{f}^{\sigma}(x_{N+1})=\sup_{\begin{subarray}{c}f\in\mathcal{H}_{k^{f}},\\ w\in\mathcal{H}_{k^{\sigma}}\\ \end{subarray}}\quad$ $f(x_{N+1})$ (4a)
$\mathrm{s.t.}\quad$ $f(x_{i})+w(x_{i})=y_{i},\>i=1,\ldots,N,$ (4b)
@@
defines the maximum norm (4c) in the RKHS $\mathcal{H}_{k^{f}+k^{\sigma}}$ based on Assumption˜1, reduced by the minimum norm required to interpolate the data.
@@
Let Assumptions˜1 and 2 hold. Then, the solution of Problem (4) is given by
$\overline{f}^{\sigma}(x_{N+1})=f^{\mu}_{\sigma}(x_{N+1})+\beta_{\sigma}\sqrt{\Sigma^{f}_{\sigma}(x_{N+1})}.$ (8)
Sketch of proof: First, following arguments from the Representer Theorem [Kimeldorf and Wahba, 1971; Schölkopf et al., 2001], we show that the solution of Problem (4) is finite-dimensional. Next, two coordinate transformations are employed to reduce the number of free variables resulting from the interpolation constraint (4b), and to address the possible rank-deficiency of the kernel matrix $K^{f}_{1:N+1,1:N+1}$ for the latent function at the test and training input locations. Finally, the problem is reduced to an equivalent linear program with a norm-ball constraint that can be analytically solved. Expressing the solution in terms of the original coordinates then leads to (8). The detailed proof can be found in Appendix˜B.
@@ 3.2 Optimal solution
Let Assumptions˜1 and 2 hold. Then, the solution of Problem (2) is given by
@@
Sketch of proof: Similarly to Lemma˜1, we first show that Problem (2) admits a finite-dimensional representation. The latter is analyzed depending on which of the constraints (2c) and (2d) are active, i.e., influence the optimal solution and have a corresponding strictly positive optimal Lagrange multiplier $\lambda^{f,\star},\lambda^{w,\star}$. This leads to three non-trivial scenarios: In Case 1, it holds that $\lambda^{f,\star}>0$, $\lambda^{w,\star}=0$ and the optimal solution of (2) can be recovered by the relaxed problem for $\sigma^{\star}\rightarrow\infty$, for which the combined RKHS-norm constraint (4c) reduces to (2c). In Case 2, $\lambda^{f,\star}=0$, $\lambda^{w,\star}>0$ and $\sigma^{\star}\rightarrow 0$ recovers the optimal solution, rendering the constraint (4c) equivalent to (2d). In Case 3, both constraints are active and the optimal noise parameter is determined by $\sigma^{\star}=\sqrt{\lambda^{f,\star}/\lambda^{w,\star}}\in(0,\infty)$, i.e., the ratio of the optimal Lagrange multipliers. The set of active constraints at the optimal solution can be determined by case distinction, based on the feasibility of the primal solutions under the respective active-constraint set. Finally, it is shown that the optimal noise parameter $\sigma^{\star}$ in all three cases minimizes (9). This is illustrated in Fig.˜1, which depicts the optimal noise parameters $\sigma^{\star}_{\sup}$$\sigma^{\star}_{\inf}$, for which the relaxed upper and lower bound, $\overline{f}^{\sigma}(x_{N+1})$ and $\underline{f}^{\sigma}(x_{N+1})$, correspond to the optimal bounds, $\overline{f}(x_{N+1})$ and $\underline{f}(x_{N+1})$, respectively. The detailed proof can be found in Appendix˜C.
@@
Figure 1: Illustrative example for Theorem˜1. The optimal upper and lower bounds (solid black) for the (unknown) latent function $f^{\mathrm{tr}}$ (dashed white) are determined by the relaxed bounds (shaded) around the GP posterior mean (dotted black) for an optimal choice of noise parameter $\sigma^{\star}_{\sup}$ (upper bound) and $\sigma^{\star}_{\inf}$ (lower bound). The three upper plots show the relaxed upper and lower bounds, $\overline{f}^{\sigma}$ and $\underline{f}^{\sigma}$ for the values $\sigma=\{10^{2},10^{0},10^{-2}\}$, respectively. The two bottom colorbars indicate the respective optimal values $\sigma^{\star}_{\sup}$ and $\sigma^{\star}_{\inf}$ for the upper and lower bound. The plotted relaxed upper (lower) bounds equal the optimal upper (lower) bound for each test point where the color of the shaded area matches the color indicated in the colorbar for the optimal value $\sigma^{\star}_{\sup}$ ($\sigma^{\star}_{\inf}$).
@@
Theorem˜1 reduces the solution of the infinite-dimensional optimization problem (2) to a scalar, unconstrained optimization problem over the noise parameter $\sigma$. As such, it is amenable for efficient iterative optimization. Since running a fixed number of iterations of, e.g., gradient descent applied to Problem (9) returns a valid, improved upper bound $\overline{f}^{\sigma}(x_{N+1})$, this allows for iterative refinement of the uncertainty envelope. The solution thereby obtained can thus be easily integrated into existing pipelines for downstream tasks, such as uncertainty quantification in streaming-data settings or model-based reinforcement learning [Deisenroth and Rasmussen, 2011; Berkenkamp et al., 2017; Kamthe and Deisenroth, 2018].
@@ 3.3 Special cases
For both cases with only one active constraint, the optimal bound can be determined directly in closed form, without optimizing for the noise parameter $\sigma$. In the following, we provide the respective optimal solutions, as well as easy-to-evaluate expressions for determining the active constraint set. Noteworthy, the analytic solutions recover known bounds in specific regression settings, highlighting that the proposed bound is a generalization thereof; we detail these connections in Section˜4.1.
@@
Let Assumptions˜1 and 2 hold. Define $P\doteq(\Phi_{1:N}^{\top}\left(K^{w}_{1:N,1:N}\right)^{-1}\Phi_{1:N})^{-1}$ and $\theta^{\mu}\doteq P\Phi_{1:N}^{\top}\left(K^{w}_{1:N,1:N}\right)^{-1}y$. Then, if
@@
Case 2 can happen in two scenarios: For finite-dimensional hypothesis spaces, i.e., $f^{\mathrm{tr}}(\cdot)=\begin{bmatrix}\phi_{1}(\cdot)&\ldots&\phi_{r}(\cdot)\end{bmatrix}\theta^{\mathrm{tr}}$ for some features $\phi_{i}(\cdot)$, $i=1,\ldots,r$, the latent function $f^{\star}(x_{N+1})$ generally does not have sufficient degrees of freedom to interpolate an arbitrary data set. As such, the optimal bound $\overline{f}(x_{N+1})$ in (13) consists of two components, the value of the least-squares estimator $f^{\mu}(x_{N+1})\doteq\Phi_{N+1}\theta^{\mu}$, as well as a term proportional to the maximum RKHS norm of the noise $\Gamma_{w}^{2}$, subtracted by $y^{\top}(K^{w}_{1:N,1:N})^{-1}y-\|\theta^{\mu}\|_{P^{-1}}$, the minimum RKHS norm required to eliminate the offset between the least-squares estimator and the data. For infinite-dimensional hypothesis spaces, the latent function $f^{\star}$ can generally interpolate the offset between the optimal noise function $w^{\star}$ and the training data; however, neglecting the RKHS-norm constraint (2c) on $f^{\star}$ only leads to sensible estimates when the test point coincides with a training input location. In this case, the feature vector $\Phi_{N+1}\in\mathbb{R}^{(N+1)\times r}$ has rank $N$, which simplifies the general result in Proposition˜2.
@@
Let Assumptions˜1 and 2 hold. Suppose that $K^{f}_{1:N,1:N}$ is invertible and $x_{N+1}=x_{k}\in\mathbb{X}$. Then, if
$\left\|\theta^{\star}\right\|_{2}^{2}$ $\doteq\left\|y+K^{w}_{1:N,k}\frac{\Gamma_{w}}{\sqrt{K^{w}_{k,k}}}\right\|_{(K^{f}_{1:N,1:N})^{-1}}^{2}\leq\Gamma_{f}^{2},$ (14)